OpenStudy (anonymous):

Two boxes of masses m and 3m are stacked. The surface between the more massive box and the horizontal surface is smooth and the surface between the boxes is rough. If the less massive box does not slide on the more massive box, what is the static friction force on the less massive box?

OpenStudy (anonymous):

Here's the picture

OpenStudy (ookawaiioo):

static friction = does not move So your only looking at m and 3m. F=usN = us*mass*gravity

OpenStudy (anonymous):

what do i have to find first ?

OpenStudy (anonymous):

The weird part about this question is that they don't tell you if the larger box is experiencing any force or not; so it's sort of an incomplete question :P

OpenStudy (anonymous):

Hmmm thanks.. let me try again

OpenStudy (anonymous):

wait did you see the picture that came with the question tho ? i posted it :o there's force moving right

OpenStudy (anonymous):

You're totally right! sorry :P

OpenStudy (anonymous):

ugh im so confused

OpenStudy (anonymous):

sorry, I meant that you're right that there is a force. It made my solution wrong, so I'm trying to work it out. Sorry for confusingness :P :/

OpenStudy (anonymous):

it's ok :D thanks for helping me

OpenStudy (anonymous):

So this is a weird one For the big box, the singular force just equals ma. \[\sum F_{big} = F = Ma\] and so it's experiencing an acceleration of \[ a = \frac{F}{M} = \frac{F}{3m}\] For the little one, it's "feeling" the big box's acceleration; so even though the static friction is keeping it from falling off the block, it's technically the force of friction that's moving it forward in the frame of the little box; so \[ \sum F_{tiny} = F_f = ma\] and the acceleration is what we found up top! \[ F_f = m\Big(\frac{F}{3m}\Big)\] So the force if friction is \[F_f = \frac{F}{3}\] Does that make sense?

OpenStudy (anonymous):

Oh.... wait

OpenStudy (anonymous):

Is that not the right answer? :/

OpenStudy (anonymous):

oHH so you basically just substitute

OpenStudy (anonymous):

i get it... thanks :D

OpenStudy (anonymous):

Yup yup ^_^ All you need to know is the acceleration of the "box system," and with that the frictional force is just what keeps the little box traveling at the same speed as the big box ^_^

OpenStudy (anonymous):

and welcome! ^^